light of wavelength 560nm passes through a double slit system that has a slit separartion of .541 mm. how far away must a screen e placed in order for the first dark fringe to appear directly oppisite both slits? units of m
Answer #1

the distance of minimum from one end of the central maximum is
Y = (2m+1)?L / 2d
y=d/2 for m=1 for 1st minima to be infront of slit

L= d2/3\lambda = 15.7 cm